# A Gaussian smooth transition vector autoregressive model: An application to the macroeconomic effects of severe weather shocks

### v2

@@
We introduce a new smooth transition vector autoregressive model with a Gaussian conditional distribution and transition weights that, for a $p$th order model, depend on the full distribution of the preceding $p$ observations. Specifically, the transition weight of each regime increases in its relative weighted likelihood. This data-driven approach facilitates capturing complex switching dynamics, enhancing the identification of gradual regime shifts. In an empirical application to the macroeconomic effects of a severe weather shock, we find that in monthly U.S. data from 1961:1 to 2022:3, the impacts of the shock are stronger in the regime prevailing in the early part of the sample and in certain crisis periods than in the regime dominating the latter part of the sample. This suggests overall adaptation of the U.S. economy to increased severe weather over time.
@@ 1 Introduction
Smooth transition vector autoregressive (STVAR) models have gained attention in the literature due to their ability to capture nonlinear dynamics in time series data (see, e.g., Hubrich and Teräsvirta, 2013, and the references therein). They provide a flexible framework for modelling autoregressive systems that exhibit gradual shifts in parameter values. Moreover, structural STVAR models allow to trace out the causal effects of economic shocks, which may be asymmetric with respect to the sign and size of the shock as well as to the initial state of the economy.
STVAR models extend the conventional vector autoregressive (VAR) models by allowing for smooth transitions between regimes, each of which is characterized by a linear vector autoregression with different autoregressive coefficients or error term covariance matrices. In this paper, we consider STVAR models in which, at each point of time, the observation is a weighted average of the conditional means of the regimes plus a random error whose covariance matrix is a weighted average of the covariance matrices of the regimes. Different STVAR models are obtained by specifying the transition weights or the error distribution in various ways. In our model, the weights are expressed in terms of time-varying transition weights that depend on the preceding observations.
@@
In particular, we introduce a new STVAR model that we call the Gaussian smooth transition vector autoregressive (GSTVAR) model. It assumes a Gaussian conditional distribution and adopts transition weights that are similar to the mixing weights of the Gaussian mixture vector autoregressive (GMVAR) model of Kalliovirta et al. (2016). For a $p$th order model, the transition weights depend on the full distribution of the preceding $p$ observations, thereby allowing to capture complicated switching dynamics. Specifically, the greater the relative weighted likelihood of a regime is, the greater its transition weight is, which facilitates associating statistical characteristics and economic interpretations to the regimes. In contrast to the GMVAR model, which incorporates unobserved discrete regime switches, our GSTVAR model allows for capturing gradual shifts in the dynamics of the data. The GSTVAR model also has the advantage in structural analysis that, unlike in the GMVAR model, identified structural shocks can be recovered from the fitted model.
@@
A variety of alternative STVAR models have been proposed in the previous literature, each characterized by a different transition weight function. A popular choice is to specify logistic transition weights and employ a logistic STVAR (LSTVAR) model (Anderson and Vahid, 1998). Another common alternative is the threshold VAR (TVAR) model (Tsay, 1998), which features discrete regimes such that the regime switches when a switching variable passes some threshold value. The LSTVAR and TVAR models have clear interpretations of the regimes and can accommodate exogenous as well as endogenous switching variables. Therefore, they are well-suited for studying problems where the transition weights are expected to depend on the level of some specific switching variables. In comparison, our GSTVAR model facilitates capturing more complicated switching dynamics and results in more data-driven regimes. However, it is not always clear how to interpret the regimes economically, and due to the complexity of the transition weights, the estimation of the GSTVAR model can be tedious in practice.
Another popular branch of the literature on nonlinear VAR models focuses on so-called Markov-switching VAR (MS-VAR) models (Krolzig, 1997, see also Hamilton, 1989, 1990). They incorporate discrete regime switches that are random and unobserved, and the switching probabilities depend on the preceding regime. The MS-VAR models can flexibly capture nonlinearities in the data, but unlike our GSTVAR model, they do not accommodate gradual shifts in the dynamics nor complicated switching dynamics.
In addition to nonlinear VAR models, time-varying parameter VAR (TVP-VAR) models (Cogley and Sargent, 2001; Cogley and Sargent, 2005; Primiceri, 2005; Koop et al., 2009; among others) allow for capturing random but persistent time-variation in parameter values. However, compared to our GSTVAR model, they have the major limitation that the stochastic processes governing the changes in parameter values are exogenous to the variables included. Consequently, changes in the macroeconomic variables have no influence on how the parameter values evolve, and, thus, potentially important endogenous dynamics of the economy are overlooked.
@@
We apply our GSTVAR model to study the macroeconomic effects of severe weather shocks and consider a monthly U.S. data set covering the period from 1961:1 to 2022:3 and containing an indicator of severe weather and a number of macroeconomic variables. We find a specification with two regimes sufficient. The first regime mainly dominates the earlier part of the sample period, particularly during the volatile periods of 1970’s and 1980’s, but also prevails in the later sample during the Financial crisis and the COVID-19 crisis. The second regime dominates the latter part of the sample period. Following Kim et al. (2022), we identify the severe weather shock recursively, placing the severe weather indicator first in the vector of variables. In both regimes, a positive severe weather shock decreases GDP, consumer prices, and the interest rate, but the effects are clearly stronger in the first regime prevailing in the earlier part of the sample period and in certain crisis periods. In contrast to Kim et al. (2022), we thus conclude that the effects of the severe weather shock are stronger in the earlier than the latter part of the sample, which suggests that the U.S. economy has adapted to the changing distribution of weather related shocks over time. However, the potential adaptation may not provide sufficient resilience when the economy is in a vulnerable state.
@@
The rest of this paper is organized as follows. Section 2 first presents our framework for STVAR models, then discusses stationarity and ergodicity of the STVAR models. We also introduce the structural STVAR model and some tools useful in empirical analysis based on it. Finally, the Gaussian STVAR (GSTVAR) model is introduced by specifying the transition weights and the error distribution, and the GSTVAR model is compared to related models proposed in the previous literature. Section 3 discusses model selection and estimation of the parameters by the method of maximum likelihood. Section 4 presents the empirical application to the macroeconomic effects of severe weather shocks. Finally, Section 5 concludes. Further discussion on the stationarity condition and a Monte Carlo algorithm for estimating the generalized impulse response function (Koop et al., 1996), as well as details on the empirical application are deferred to appendices. The introduced methods are implemented to the accompanying R package sstvars (Virolainen, 2024), available in the CRAN repository.
@@ 2.1 General framework for STVAR models
where $\phi_{1,0},...,\phi_{M,0}\in\mathbb{R}^{d}$, $m=1,...,M$, are the intercept parameters, $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregressive matrices, and $\Omega_{1},...,\Omega_{M}$ are the positive definite $(d\times d)$ covariance matrices of the regimes. The serially uncorrelated reduced form innovations $u_{t}$ follow some distribution with zero mean and conditional covariance matrix $\Omega_{y,t}$. The transition weights $\alpha_{m,t}$ are assumed to be $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$ and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. These weights convey the relative proportions of the regimes at each point in time and determine how the process shifts between them. Specifically, when the process is completely in one of the regimes, the transition weight $\alpha_{m,t}$ of that regime equals unity, while the weights of the rest of the regimes are equal to zero. As the process begins a shift towards another regime, the transition weight of the emerging regime increases due to the dynamics captured in the transition weight function through the preceding observations. Concurrently, the weight of the previously dominant regime decreases.
@@
We assume that the transition weights are functions of $\{y_{t-j},j=1,...,p\}$, which is required for applicability of the stationarity condition discussed below. We also assume that the transition weights are identical for all the individual equations in (2.1) and (2.3), which is also required for applicability of the stationarity condition. Consequently, at each $t$‚ the process can be described as a weighted sum of linear vector autoregressions. As discussed in Section 2.3.1, different STVAR models are obtained by specifying the transition weights or the error distribution in various ways. For instance, defining $\alpha_{2,t}$ as the logistic function and $\alpha_{1,t}$ as $1-\alpha_{2,t}$ gives the two-regime logistic STVAR model popular in the empirical macroeconomic literature (see, e.g., Auerbach and Gorodnichenko, 2012), while transition weights like the mixing weights in Kalliovirta et al. (2016) yield the Gaussian STVAR model discussed in Section 2.3.
For standard asymptotic inference, the STVAR process must be stationary and ergodic. To verify that this is indeed the case, we rely on the sufficient condition for stationarity and ergodicity derived by Kheifets and Saikkonen (2020). However, their parametrization is not quite the same as ours, so their result must be slightly modified for our purposes (see Appendix A for details). In terms of the companion form AR matrices of the regimes defined as
@@
$\rho(\mathcal{A})=\underset{j\rightarrow\infty}{\limsup}\left(\underset{A\in\mathcal{A}^{j}}{\sup}\rho(A)\right)^{1/j},$ (2.5)
@@
It is worth to mention that a necessary condition for the sufficient condition is that the usual stability condition is satisfied for each of the regimes since $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein). Therefore, the following condition, which is analogous to Corollary 1 of Kheifets and Saikkonen (2020), is necessary for Condition 1.
@@
Since the sufficient Condition 1 is computationally costly to verify with a reasonable accuracy (see, e.g., Chang and Blondel, 2013), it is useful to employ the more easily verified necessary Condition 2 in numerical estimation discussed in Section 3. As our estimation procedure produces a set of alternative local solutions, Condition 1 can be checked for each of them after the estimation, and the researcher can choose to use the best local solution whose stationarity can be verified.
A number of methods for bounding the JSR have been proposed in the literature, many of which are discussed by Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2024) implements the branch-and-bound method of Gripenberg (1996). The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods in the estimation of the JSR to enhance computational efficiency.
@@ 2.2 Structural STVAR model
To conduct structural analysis in the STVAR model, we need to find its structural counterpart involving orthogonal, serially uncorrelated structural errors, or shocks, $e_{t}$. This amounts to finding a non-singular ($d\times d)$ impact matrix $B_{t}$ such that the conditional covariance matrix of $e_{t}=B_{t}^{-1}u_{t}$, conditional on $\mathcal{F}_{t-1}$, is a diagonal matrix (typically normalized to the identity matrix). In other words, $B_{t}$ must be such that $B_{t}^{-1}\Omega_{y,t}B_{t}^{\prime-1}=I_{d}$. However, without additional identifying restrictions $B_{t}$ satisfying this equation is not unique, and as discussed by, e.g., Kilian and Lütkepohl (2017, Chapter 18), imposing such restrictions in nonlinear structural VAR models may not be straightforward. In our empirical application in Section 4, we consider a recursive model, where $B_{t}$ is obtained by a lower-triangular Cholesky decomposition of the covariance matrix of the reduced-form residuals in each regime.
Conventional impulse responses can be computed separately for each regime in the structural STVAR model, as has been done in much of the nonlinear structural VAR literature. However, by thus precluding future regime switches, this approach changes the structure of the model and makes the impulse response analysis subject to Lucas critique. Moreover, the expected effects of the structural shocks generally depend on the initial values of the variables as well as the sign and size of the shock, which is not accounted for by the conventional impulse responses. However, the generalized impulse reponse function (GIRF) of Koop et al. (1996) can be used. The GIRF is defined as
@@
where $h$ is the chosen horizon and $\mathcal{F}_{t-1}=\sigma\{y_{t-j},j>0\}$ as before. The first term on the right side of (2.6) is the expected realization of the process at time $t+h$ conditionally on a structural shock of sign and size $\delta_{j}\in\mathbb{R}$ in the $j$th element of $e_{t}$ at time $t$ and the previous observations. The latter term on the right side is the expected realization of the process conditionally on the previous observations only. The GIRF thus expresses the expected difference in the future outcomes when the structural shock of sign and size $\delta_{j}$ in the $j$th element hits the system at time $t$ as opposed to all shocks being random. An interesting feature of the structural STVAR model is that besides the generalized impulse response functions of the observable variables, it produces the GIRFs of the transition weights $\alpha_{m,t}$, $m=1,...,M$, which yield information on the dynamic effects of the shocks on regime switches. They are obtained by replacing $y_{t+h}$ with $\alpha_{m,t}$ on the right side of Equation (2.6).
It is easy to see that the GSTVAR model has a $p$-step Markov property, so conditioning on (the $\sigma$-algebra generated by) the $p$ previous observations $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ is effectively the same as conditioning on $\mathcal{F}_{t-1}$ at time $t$ and later. In order to estimate GIRFs conditional on that the economy is in a specific regime when the shock arrives, we generate initial values $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ from the stationary distribution of that regime and the estimate GIRFs for each of these initial values. Taking the sample mean and quantiles over the estimated GIRFs then produces the point estimate and confidence bounds, respectively. Our confidence bounds thus reflect uncertainty about the initial state of the economy within the specific regime. A more detailed description of our Monte Carlo algorithm for estimating the GIRF is presented in Appendix B.

### v3

@@
We introduce a new smooth transition vector autoregressive model with a Gaussian conditional distribution and transition weights that, for a $p$th order model, depend on the full distribution of the preceding $p$ observations. Specifically, the transition weight of each regime increases in its relative weighted likelihood. This data-driven approach facilitates capturing complex switching dynamics, enhancing the identification of gradual regime shifts. In an empirical application to the macroeconomic effects of a severe weather shock, we find that in monthly U.S. data from 1961:1 to 2022:3, the shock has stronger impact in the regime prevailing in the early part of the sample and in certain crisis periods than in the regime dominating the latter part of the sample. This suggests overall adaptation of the U.S. economy to severe weather over time.
@@ 1 Introduction
Smooth transition vector autoregressive (STVAR) models that allow for gradual shifts in parameter values have become popular due to their ability to capture nonlinear dynamics in time series data (for a survey of the literature, see, e.g., Hubrich and Teräsvirta, 2013, and the references therein). They extend the conventional linear vector autoregressive (VAR) model by allowing for smooth transitions between regimes, each of which is characterized by a linear vector autoregression with different coefficients or error covariance matrices. Each observation is the sum of a weighted average of the conditional means of the regimes and a random error, whose covariance matrix is a weighted average of the covariance matrices of the regimes. Like other nonlinear structural VAR models, the structural counterparts of STVAR models enable tracing out the causal effects of economic shocks, which may depend on the initial state of the economy or on the sign or size of the shock.
@@
Different STVAR models are obtained by specifying the transition weights or the error distribution in various ways. In this paper, we introduce the Gaussian smooth transition vector autoregressive (GSTVAR) model, where the transition weights are similar to the mixing weights of the Gaussian mixture vector autoregressive (GMVAR) model of Kalliovirta et al. (2016). For a $p$th order model, the transition weights in both models depend on the full distribution of the preceding $p$ observations, which enables capturing complicated switching dynamics. Specifically, the greater the relative weighted likelihood of a regime is, the greater its transition weight is, which facilitates associating statistical characteristics and economic interpretations to the regimes. However, in contrast to the GMVAR model, which incorporates unobserved discrete regime switches, our GSTVAR model has the advantage that it allows for capturing gradual shifts in the dynamics of the data. Moreover, the identified shocks of the structural GSTVAR model can be recovered from the fitted model and used in impulse response analysis to further reflect the properties of the data, which is not the case with the structural GMVAR model.
@@
Probably the most popular among the STVAR models put forward in the previous literature is the logistic STVAR (LSTVAR) model (Anderson and Vahid, 1998) characterized by logistic transition weights. In the LSTVAR model, the interpretation of the regimes is clear, and it can accommodate exogenous as well as endogenous switching variables. These properties are shared by the often employed threshold VAR (TVAR) model (Tsay, 1998), albeit the latter features discrete regimes instead of allowing for smooth transitions between them. Hence, these models are well-suited, when the transition weights can be expected to depend on the level of some specific switching variables. In contrast, our GSTVAR model facilitates capturing more complicated switching dynamics and results in more clearly data-driven regimes, but it may not always be straightforward to interpret the regimes economically.
Besides STVAR and TVAR models, there are a number of alternative approaches to modeling nonlinearities in multivariate time series, including so-called Markov-switching VAR (MS-VAR) models (Krolzig, 1997, see also Hamilton, 1989, 1990) and time-varying parameter VAR (TVP-VAR) models (Cogley and Sargent, 2001; Cogley and Sargent, 2005; Primiceri, 2005; Koop et al., 2009; among others). The former incorporate discrete regime switches that are random and unobserved, and the switching probabilities depend on only the preceding regime. While they can flexibly capture nonlinearities in the data, unlike our GSTVAR model, they do not accommodate gradual shifts in the dynamics nor complicated switching dynamics. As to TVP-VAR models, compared to our GSTVAR model, they have the major limitation that the stochastic processes governing the changes in parameter values are exogenous to the variables included, which may lead to overlooking important endogenous dynamics.
We apply the structural GSTVAR model to find out about the macroeconomic effects of severe weather. This issue was recently studied by Kim et al. (in press) by a STVAR model with transition weights defined by a linear time trend, which only allows a single switch from one regime to the other. Following them, we include an indicator of severe weather in addition to a number of macroeconomic variables in the model and estimate it on monthly U.S. data from 1961:1 to 2022:3. In line with Kim et al. (in press), we find evidence in favor of nonlinearity, and a GSTVAR model with two regimes is deemed sufficient. However, our regimes are characterized by features quite different from theirs. One regime dominates the earlier part of the sample period, particularly the volatile periods of the 1970s and 1980s, and also prevails later during the Financial crisis and the COVID-19 crisis, while the latter part of the sample period is predominantly characterized by the other regime.
@@
In both regimes, a recursively identified positive severe weather shock decreases GDP, consumer prices, and the interest rate, but the effects are clearly stronger in the regime prevailing in the earlier part of the sample period and in certain crisis periods. In contrast to Kim et al. (in press), who found the effects of severe weather stronger in the latter part of the sample period, we thus conclude that the U.S. economy has adapted to the changing distribution of weather related shocks over time. However, since strong effects are found also in the crisis periods, it seems that the apparent adaptation does not provide sufficient resilience when the economy is in a vulnerable state.
@@
The rest of this paper is organized as follows. Section 2 first presents our framework for STVAR models and then discusses their stationarity and ergodicity. We also introduce the structural STVAR model and some tools useful in empirical analysis based on it. Finally, the new Gaussian STVAR (GSTVAR) model is introduced by specifying the transition weights and the error distribution. In Section 3, model selection and estimation of the parameters by the method of maximum likelihood are discussed. Section 4 contains the empirical application to the macroeconomic effects of severe weather shocks. Section 5 concludes the paper. Further details can be found the in the appendices, and the introduced methods are implemented to the accompanying R package sstvars (Virolainen, 2025a), available in the CRAN repository.
@@ 2.1 General framework for STVAR models
where $\phi_{1,0},...,\phi_{M,0}\in\mathbb{R}^{d}$, $m=1,...,M$, are the intercept parameters, $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregressive matrices, and $\Omega_{1},...,\Omega_{M}$ are the positive definite $(d\times d)$ covariance matrices of the regimes. The serially uncorrelated reduced form innovations $u_{t}$ follow some distribution with zero mean and conditional covariance matrix $\Omega_{y,t}$.
The transition weights $\alpha_{m,t}$ are assumed to be $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$ and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. These weights convey the relative proportions of the regimes at each point in time and determine how the process shifts between them. Specifically, when the process is completely in one of the regimes, the transition weight $\alpha_{m,t}$ of that regime equals unity, while the weights of the rest of the regimes are equal to zero. As the process begins a shift towards another regime, the transition weight of the emerging regime increases due to the dynamics captured in the transition weight function through the preceding observations. Concurrently, the weight of the previously dominant regime decreases.
@@
For standard asymptotic inference, the STVAR process must be stationary and ergodic. To verify that this is indeed the case, we rely on the sufficient condition for stationarity and ergodicity derived by Kheifets and Saikkonen (2020) (based on the more general results of Saikkonen, 2008). However, their parametrization is not quite the same as ours, so their result must be slightly modified for our purposes (see Appendix A for details). In terms of the companion form AR matrices of the regimes defined as
@@
$\rho(\mathcal{A})=\underset{j\rightarrow\infty}{\limsup}\left(\underset{A\in\mathcal{A}^{j}}{\sup}\rho(A)\right)^{1/j}$ (2.5)
@@
It is worth mentioning that a necessary condition for Condition 1 is that the usual stability condition is satisfied for each of the regimes, as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein). Therefore, the following condition, which is analogous to Corollary 1 of Kheifets and Saikkonen (2020), is necessary for Condition 1.
@@
Since the sufficient Condition 1 is computationally costly to verify with reasonable accuracy (see, e.g., Chang and Blondel, 2013), it is useful to employ the more easily verified necessary Condition 2 in numerical estimation (see Section 3). As our estimation procedure produces a set of alternative local solutions, Condition 1 can be checked for each of them after the estimation, and the best local solution is selected among those for which stationarity can be verified.
A number of methods for bounding the JSR have been proposed in the literature, many of which are discussed by Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996). The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods in the estimation of the JSR to enhance computational efficiency.
@@ 2.2 Structural STVAR model
To conduct structural analysis, we need to find the structural counterpart of the STVAR model that involves orthogonal, serially uncorrelated structural errors, or shocks, $e_{t}$. This amounts to finding a non-singular ($d\times d)$ impact matrix $B_{t}$ such that the conditional covariance matrix of $e_{t}=B_{t}^{-1}u_{t}$, conditional on $\mathcal{F}_{t-1}$, is a diagonal matrix (typically normalized to the identity matrix). In other words, $B_{t}$ must be such that $B_{t}^{-1}\Omega_{y,t}B_{t}^{\prime-1}=I_{d}$. However, without additional identifying restrictions $B_{t}$ satisfying this equation is not unique, and as discussed by, e.g., Kilian and Lütkepohl (2017, Chapter 18), imposing such restrictions in nonlinear structural VAR models may not be straightforward. In our empirical application in Section 4, we consider a recursive model, where $B_{t}$ is obtained by a lower-triangular Cholesky decomposition of the conditional covariance matrix $\Omega_{y,t}$.
Conventional impulse responses can be computed separately for each regime in the structural STVAR model, as has been done in much of the nonlinear structural VAR literature. However, by thus precluding future regime switches, this approach changes the structure of the model and makes the impulse response analysis subject to Lucas critique. Moreover, the expected effects of the structural shocks generally depend on the initial values of the variables as well as on the sign and size of the shock, which is not accounted for by the conventional impulse responses. An appropriate alternative is the generalized impulse response function (GIRF) of Koop et al. (1996), defined as
@@
where $h$ is the horizon and $\mathcal{F}_{t-1}=\sigma\{y_{t-j},j>0\}$ as before. The first term on the right side of (2.6) is the expected realization of the process at time $t+h$ conditional on a structural shock of sign and size $\delta_{j}\in\mathbb{R}$ in the $j$th element of $e_{t}$ at time $t$ and the previous observations. The latter term on the right side is the expected realization of the process conditional on the previous observations only. The GIRF thus expresses the expected difference in the future outcomes when the structural shock of sign and size $\delta_{j}$ in the $j$th element hits the system at time $t$ as opposed to all shocks being random. An interesting feature of the structural STVAR model is that besides the generalized impulse response functions of the observable variables, it produces the GIRFs of the transition weights $\alpha_{m,t}$, $m=1,...,M$, which yield information on the dynamic effects of the shocks on regime switches. They are obtained by replacing $y_{t+h}$ with $\alpha_{m,t}$ on the right side of Equation (2.6).
Typically STVAR models have a $p$-step Markov property, and this is also the case with our GSTVAR model to be introduced in Section 2.3. Thus conditioning on (the $\sigma$-algebra generated by) the $p$ previous observations $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ is effectively the same as conditioning on $\mathcal{F}_{t-1}$ at time $t$ and later. To estimate the GIRFs conditional on the economy being in a specific regime, say $\tilde{m}$, when the shock arrives, we select the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data that indicate the dominance of this regime in the time period $t$. That is, we select the histories $\boldsymbol{y}_{t-1}$ for which the transition weight $\alpha_{\tilde{m},t}$ is large, say, larger than $0.75$, indicating that the process is largely in Regime $\tilde{m}$. For each of these histories $\boldsymbol{y}_{t-1}$, we compute the GIRFs to the structural shock $e_{it}$ recovered from the data (the $i$th element of the estimate of $B_{t}^{-1}(y_{t}-\mu_{y,t})$), so they reflect the properties of the data. To then make GIRFs to shocks with different signs and sizes comparable, they are all scaled to correspond to some fixed (say, unit) instantaneous increase in one of the variables. Finally, the distribution of GIRFs obtained for all histories $\boldsymbol{y}_{t-1}$ can be superimposed in a so-called “shotgun plot” (cf. Inoue and Kilian, 2016). When a low level of opacity is used, the darkness of each region in the figure indicates the concentration of GIRFs in it (for an example, see Figure 2). A detailed description of our Monte Carlo algorithm for estimating the GIRF is presented in Appendix B. [^1] Another way to estimate GIRFs conditional on the economy being in a specific regime when the shock arrives is to generate the initial values $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ from the stationary distribution of that regime. Comparison of regime-dependent GIRFs for different signs and sizes of the shock then reveals asymmetries with respect to these features within the regime.